# Anabelian aspects of the outer automorphism groups of the absolute Galois groups of mixed-characteristic local fields

### v1

@@ 5. $(d)$
Next, we verify assertion (2). Since $\mathcal{O}_{k^{(d=1)}}^{\times}$ is a profinite group, to show that the subgroup $\varphi^{\times}(\mathcal{O}_{k^{(d=1)}}^{\times})\cap\mathcal{O}_{k^{(d=1)}}^{\times}\subset\mathcal{O}_{k^{(d=1)}}^{\times}$ is not open in $\mathcal{O}_{k^{(d=1)}}^{\times}$, it suffices to show that it is not a subgroup of finite index. Moreover, it follows from [Hsh3], Lemma 1.2, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$, and the definition of perfection [cf. §0, Notational conventions, Monoids] that it suffices to show $\varphi_{+}(k^{(d=1)}_{+})\cap k^{(d=1)}_{+}=0$ in order to prove that the subgroup
@@
Next, we show that $\rho$ is a crystalline representation. Let $F_{\pi}$ be the Lubin-Tate formal group law associated with the Lubin-Tate formal power series $\pi t+t^{{p_{k}}^{f_{k}}}$. Then $\rho$ is isomorphic to the continuous $p_{k}$-adic representation induced by the action on the $p_{k}$-adic Tate module of $F_{\pi}$ [cf., e.g., [Ser], $\rm I\hskip-1.49994ptI\hskip-1.49994ptI$, §A.4, Proposition 4]. Since $F_{\pi}$ is defined over $\mathcal{O}_{k}$, the corresponding $p_{k}$-divisible group is also defined over $\mathcal{O}_{k}$. It is well-known [cf. [Fon], Theorem 6.2, $\rm\hskip 1.79993pti\hskip 1.79993pt)$] that the continuous $p_{k}$-adic representation induced by the action on the $p_{k}$-adic Tate module of $F_{\pi}$ is crystalline. Thus, the representation $\rho$ is a crystalline representation.
@@
is an automorphism that restricts to an automorphism of the subgroup $\mathcal{O}_{k^{(d=1)}}^{\times}\subset\mathcal{O}_{k}^{\times}$. We write $\varphi^{\operatorname{ab}}\colon G_{k}^{\operatorname{ab}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}G_{k}^{\operatorname{ab}}$ for the automorphism of the topological group $G_{k}^{\operatorname{ab}}$ induced by $\varphi$. It follows from the assumption of the contradiction that $\rho\circ\varphi$ is Hodge–Tate. Thus, it follows from [Hsh4], Lemma 1.9, that there exists an open subgroup $U\subset\mathcal{O}_{k^{(d=1)}}^{\times}$ whose image by the composite
@@
Suppose that $d_{k}=1$ and that $p_{k}$ is odd. As observed in the proof of Theorem 1.9, there exists an irreducible, abelian, and crystalline [hence, Hodge-Tate] one-dimensional continuous representation $\rho_{0}\colon G_{k}\to\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(V)$. As shown in [Hsh4], Theorem A, when a continuous $p_{k}$-adic representation $\rho\colon G_{k}\to\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(V)$ is one-dimensional, or, even, two-dimensional and reducible, the property that $\rho$ is Hodge-Tate is equivalent to the property that $\rho$ is Aut-intrinsically Hodge-Tate. Thus, this representation $\rho_{0}\colon G_{k}\to\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(V)$ is Aut-intrinsically Hodge-Tate.
On the other hand, it is well-known [cf., e.g., the discussion given at the final portion of [N-S-W], Chapter VII, §5] that there exists an outer automorphism of $G_{k}$ that is not induced by any field automorphism of $k$. Thus, it follows from [Hsh1], Corollary 3.4, that there exists a continuous representation of $G_{k}$ that is Hodge-Tate but not Aut-intrinsically Hodge-Tate. It follows that the following inclusion holds:
@@
We note that any one-dimensional continuous linear representation that is Hodge-Tate is necessarily Aut-intrinsically Hodge-Tate [cf. [Hsh4], Theorem A]. In particular, the representation $\rho^{\prime}$ is Aut-intrinsically Hodge-Tate but not crystalline. Thus, there is no inclusion relation in either direction between the classes of Aut-intrinsically Hodge-Tate representations and crystalline representations.
@@
2 On the non-normality of the field-theoretic subgroups of the outer automorphism groups of the absolute Galois groups of absolutely Galois MLFs
In the present §2, let $k$ be an MLF, $\overline{k}$ an algebraic closure of $k$, and $G$ a group of MLF-type. In the present §2, we discuss the outer automorphism groups of the absolute Galois groups of certain MLFs. In particular, we show that if $p_{k}$ is odd, $d_{k}$ is greater than one, and $k$ is an absolutely Galois MLF, i.e., the extension $k/k^{(d=1)}$ is a Galois extension, then the field-theoretic subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is not normal. This result generalizes certain theorems in [H-N] and [Hsh2].
@@
We shall refer to the image of the natural injective group homomorphism $\operatorname{Aut}(k)\hookrightarrow\operatorname{Out}(G_{k})$ [cf., e.g., [Hsh3], Proposition 2.1] as the field-theoretic subgroup of $\operatorname{Out}(G_{k})$.
@@
It is well-known that the isomorphism class of the field $k$ itself cannot be reconstructed group-theoretically from $G_{k}$ in general. For example, let $p$ be an odd prime number, $k_{0}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{Q}_{p}(\zeta_{p},\sqrt[p]{p})$ and $k_{1}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{Q}_{p}(\zeta_{p},\sqrt[p]{p+1})$, where $\zeta_{p}$ is a primitive $p$-th root of unity. Fix an algebraic closure $\overline{k_{0}}$ of $k_{0}$ and an algebraic closure $\overline{k_{1}}$ of $k_{1}$. Then it follows from the main theorem of [J-R] that the group $G_{k_{0}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Gal}(\overline{k_{0}}/k_{0})$ is isomorphic to the group $G_{k_{1}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Gal}(\overline{k_{1}}/k_{1})$, while the field $k_{0}$ is not isomorphic to the field $k_{1}$.
Furthermore, the following observation illustrates, in a stronger sense, that a group isomorphism $G_{k_{0}}\cong G_{k_{1}}$ is not compatible with their ring structures. Let $\varphi\colon G_{k_{1}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}G_{k_{0}}$ be an arbitrary group isomorphism, and let $\Phi\colon\operatorname{Out}(G_{k_{1}})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\operatorname{Out}(G_{k_{0}})$ denote the induced group isomorphism. Then it follows from [Hsh2], Corollary 8.6, that $\Phi(\operatorname{Gal}(k_{1}/\mathbb{Q}_{p}(\zeta_{p})))\not\subset N_{\operatorname{Out}(G_{k_{0}})}(\operatorname{Aut}(k_{0}))$. In particular, the subgroup $\operatorname{Aut}(k_{0})\subset\operatorname{Out}(G_{k_{0}})$ is not a normal subgroup of $\operatorname{Out}(G_{k_{0}})$.
@@
A group of MLF-type is an example of a slim profinite group [cf., e.g., [Hsh3], Proposition 1.8].
@@
Assertion (1) follows directly from Lemma 2.4, (1), and [Hsh3], Lemma 1.8. Assertion (2) follows directly from [Hsh3], Lemma 1.8. Assertion (3) follows directly from Lemma 2.4, (3), [Hsh3], Lemma 1.8, and the commutativity of the following diagram:
@@
Let $k(G_{k})$ be the topological field defined in [Hsh2], Definition 5.2 [Note that if $d_{k}=1$, then $d(G_{k})=1$ — cf. [Hsh3], Proposition 3.6]. By definition, the group $k(G_{k})^{\times}$ is the multiplicative group of the field $k(G_{k})$. Moreover, it follows immediately from [Hsh2], Theorem 5.4, that $k(G_{k})$ is canonically isomorphic to $k$ as a topological field. It follows from the functoriality of the mono-anabelian reconstruction algorithm that the following diagram commutes:
@@
This assertion follows immediately from Proposition 2.5, (3), Lemma 2.7, the commutative diagram appearing in the statement of [Hsh3], Lemma 1.7, $\rm(i\hskip-0.80002pti\hskip-0.80002pti)$, and the functoriality of the transfer map. ∎
The second main theorem of the present paper — namely, the non-normality of the field-theoretic subgroup of $\operatorname{Out}(G_{k})$ — follows from some results in §1 and the techniques developed in [H-N] under the assumptions that $d_{k}$ is even, $p_{k}$ is odd, and $k$ is absolutely Galois.
@@
Theorem 2.9 is a generalization of the main theorem of [H-N] and one of the main results of [Hsh2], which appeared in Remark 2.2. The method of the proof of Theorem 2.9 is essentially the same as the method of the proof that was applied in [H-N] and is substantially different from the approach taken in [Hsh2].
@@
Suppose that the subgroup $G_{k}\subset G_{k^{(d=1)}}$ is a characteristic open subgroup. For example, if $k/k^{(d=1)}$ is an abelian extension, then the subgroup $G_{k}\subset G_{k^{(d=1)}}$ is characteristic and open [cf. [Hsh2], Theorem 6.3, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$]. Then, by Proposition 2.5, (2), (3), we obtain the following exact sequence of groups:
@@
From the existence of this exact sequence and the fact that $\operatorname{Out}(G_{k^{(d=1)}})\neq 1$ for odd residue characteristic $p_{k}$ [cf., e.g., the discussion given at the final portion of [N-S-W], Chapter VII, §5], we deduce that $\operatorname{Aut}(k)\subsetneq N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$. Put another way, there exists an element $\alpha\in N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$ such that $\alpha\notin\operatorname{Aut}(k)$. As discussed in the discussion following Remark 2.2, in general, the subgroup $(k^{(d=1)})^{\times}\subset k^{\times}$ cannot be reconstructed from the group $G_{k}$ in a purely group-theoretic and functorial manner. Put another way, the subgroup $(k^{(d=1)})^{\times}\subset k^{\times}$ depends on the ring structure of $k$ in an essential way. On the other hand, by Theorem 2.8, if $k$ is an absolutely Galois MLF and $\alpha\in N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$, then the induced isomorphism $\alpha^{\times}$ preserves the subgroup $(k^{(d=1)})^{\times}\subset k^{\times}$ and $\alpha^{\times}|_{(k^{(d=1)})^{\times}}=\operatorname{id}_{(k^{(d=1)})^{\times}}$. It follows from this discussion that every element $\alpha\in N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$, which is not necessarily contained in $\operatorname{Aut}(k)$, preserves certain structures of ring-theoretic nature.
@@
Let $\alpha$ be an element of $N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$. Suppose that $k$ is an absolutely Galois MLF. As is shown in [H-N], Lemma 3.3, it holds that $\alpha_{+}|_{k^{(d=1)}_{+}}=\operatorname{id}_{k^{(d=1)}_{+}}$.
In what follows, we provide an alternative proof of this fact, which is stated in [H-N], by means of Theorem 2.8. What we need to show is that, for $\alpha\in\operatorname{Out}(G_{k})$, if $\alpha$ is $\mathbb{Q}_{p_{k}}^{\times}$-preserving, then it is also $(\mathbb{Q}_{p_{k}})_{+}$-preserving. Since $\alpha$ is $\mathbb{Q}_{p_{k}}^{\times}$-preserving, we have $\alpha^{\times}|_{(k^{(d=1)})^{\times}}=\operatorname{id}_{(k^{(d=1)})^{\times}}$. In particular, we have $\alpha^{\times}|_{\mathcal{O}^{\times}_{k^{(d=1)}}}=\operatorname{id}_{\mathcal{O}^{\times}_{k^{(d=1)}}}$. Therefore, on $2p_{k}\mathcal{I}_{k}$, we have $\alpha_{+}=\operatorname{id}_{2p_{k}\mathcal{I}_{k}}$. Since, for an arbitrary $x\in k_{+}$, there exists an integer $n$ such that $p_{k}^{n}x\in 2p_{k}\mathcal{I}_{k}$ [cf. [Hsh3], Lemma 1.2, $\rm(\hskip-0.59998ptv\hskip-0.80002pti)$], we conclude that $\alpha_{+}=\operatorname{id}_{k_{+}}$.
@@
Since $G$ is a topologically finitely generated profinite group [cf. Theorem 1.1], for any integer $n>0$, there exist only finitely many subgroups of $G$ of index $n$. In particular, the set of all characteristic subgroups of $G$ of finite index forms a fundamental system of open neighborhoods of $1\in G$. Hence, it follows from [R-Z], Proposition 4.4.3, that $\operatorname{Aut}(G)$ and $\operatorname{Out}(G)$ have natural profinite group structures. In the remainder of the present paper, we endow $\operatorname{Aut}(G)$ and $\operatorname{Out}(G)$ with these profinite group structures.
Let us consider the group homomorphism $\operatorname{Out}(G)\to\operatorname{Aut}_{\mathbb{Q}_{p(G)}}(k_{+}(G))$ defined via the mono-anabelian reconstruction algorithm [cf. [Hsh3], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$, [H-N], Lemma 1.2]. We equip $\operatorname{Aut}_{\mathbb{Q}_{p(G)}}(k_{+}(G))$ with the topology induced by the $p(G)$-adic topology of $k_{+}(G)$, i.e., the relative topology with respect to
@@
By construction, the map $\operatorname{Aut}(G)\to\operatorname{Aut}(\mathcal{O}^{\times}(G))$ factors through the composite $\operatorname{Aut}(G)\to\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})\to\operatorname{Aut}(\mathcal{O}^{\times}(G))$, where we write $\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})$ for the subgroup of $\operatorname{Aut}(G^{\mathrm{ab}})$ consisting of elements that preserve the subgroup $\mathcal{O}^{\times}(G)\subset G^{\mathrm{ab}}$. We equip $\operatorname{Aut}(G^{\mathrm{ab}})$ with the profinite topology as in the case of $\operatorname{Aut}(G)$. The subgroup $\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})$ is endowed with the subspace topology induced from $\operatorname{Aut}(G^{\mathrm{ab}})$. It is clear that the map $\operatorname{Aut}(G)\to\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})$ is continuous. Moreover, since the profinite topology and the compact-open topology on $\mathrm{Aut}(G^{\mathrm{ab}})$ coincide [cf., e.g., [R-Z], Theorem 4.4.2], the map $\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})\to\operatorname{Aut}(\mathcal{O}^{\times}(G))$, which is defined by $\varphi\mapsto\varphi|_{\mathcal{O}^{\times}(G)}$, is continuous. Thus, the action $\operatorname{Aut}(G)\curvearrowright\mathcal{O}^{\times}(G)$ is continuous. This completes the proof of Lemma 2.13. ∎
When $d_{k}>1$, let us consider the basis $y_{1},\ldots,y_{d_{k}}$ of $k_{+}$ over $\mathbb{Q}_{p_{k}}$ corresponding to the basis of $k_{+}(G_{k})$ that appeared in the discussion preceding Lemma 1.2 via the isomorphism of topological groups $k_{+}(G_{k})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}k_{+}$ of [Hsh3], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$. In the remainder of the present paper, we always equip $k_{+}$ with this basis, which allows us to identify $\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(k_{+})$ with $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$. We equip $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ with its natural $p_{k}$-adic Lie group structure.

### v2

@@ 5. $(d)$
Next, we verify assertion (2). Since $\mathcal{O}_{k^{(d=1)}}^{\times}$ is a profinite group, to show that the subgroup $\varphi^{\times}(\mathcal{O}_{k^{(d=1)}}^{\times})\cap\mathcal{O}_{k^{(d=1)}}^{\times}\subset\mathcal{O}_{k^{(d=1)}}^{\times}$ is not open in $\mathcal{O}_{k^{(d=1)}}^{\times}$, it suffices to show that it is not a subgroup of finite index. Moreover, it follows from [8], Lemma 1.2, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$, and the definition of perfection [cf. §0, Notational conventions, Monoids] that it suffices to show $\varphi_{+}(k^{(d=1)}_{+})\cap k^{(d=1)}_{+}=0$ in order to prove that the subgroup
@@
Next, we show that $\rho$ is a crystalline representation. Let $F_{\pi}$ be the Lubin-Tate formal group law associated with the Lubin-Tate formal power series $\pi t+t^{{p_{k}}^{f_{k}}}$. Then $\rho$ is isomorphic to the continuous $p_{k}$-adic representation induced by the action on the $p_{k}$-adic Tate module of $F_{\pi}$ [cf., e.g., [19], $\rm I\hskip-1.49994ptI\hskip-1.49994ptI$, §A.4, Proposition 4]. Since $F_{\pi}$ is defined over $\mathcal{O}_{k}$, the corresponding $p_{k}$-divisible group is also defined over $\mathcal{O}_{k}$. It is well-known [cf. [3], Theorem 6.2, $\rm\hskip 1.79993pti\hskip 1.79993pt)$] that the continuous $p_{k}$-adic representation induced by the action on the $p_{k}$-adic Tate module of $F_{\pi}$ is crystalline. Thus, the representation $\rho$ is a crystalline representation.
@@
is an automorphism that restricts to an automorphism of the subgroup $\mathcal{O}_{k^{(d=1)}}^{\times}\subset\mathcal{O}_{k}^{\times}$. We write $\varphi^{\operatorname{ab}}\colon G_{k}^{\operatorname{ab}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}G_{k}^{\operatorname{ab}}$ for the automorphism of the topological group $G_{k}^{\operatorname{ab}}$ induced by $\varphi$. It follows from the assumption of the contradiction that $\rho\circ\varphi$ is Hodge–Tate. Thus, it follows from [9], Lemma 1.9, that there exists an open subgroup $U\subset\mathcal{O}_{k^{(d=1)}}^{\times}$ whose image by the composite
@@
Suppose that $d_{k}=1$ and that $p_{k}$ is odd. As observed in the proof of Theorem 1.9, there exists an irreducible, abelian, and crystalline [hence, Hodge-Tate] one-dimensional continuous representation $\rho_{0}\colon G_{k}\to\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(V)$. As shown in [9], Theorem A, when a continuous $p_{k}$-adic representation $\rho\colon G_{k}\to\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(V)$ is one-dimensional, or, even, two-dimensional and reducible, the property that $\rho$ is Hodge-Tate is equivalent to the property that $\rho$ is Aut-intrinsically Hodge-Tate. Thus, this representation $\rho_{0}\colon G_{k}\to\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(V)$ is Aut-intrinsically Hodge-Tate.
On the other hand, it is well-known [cf., e.g., the discussion given at the final portion of [15], Chapter VII, §5] that there exists an outer automorphism of $G_{k}$ that is not induced by any field automorphism of $k$. Thus, it follows from [6], Corollary 3.4, that there exists a continuous representation of $G_{k}$ that is Hodge-Tate but not Aut-intrinsically Hodge-Tate. It follows that the following inclusion holds:
@@
We note that any one-dimensional continuous linear representation that is Hodge-Tate is necessarily Aut-intrinsically Hodge-Tate [cf. [9], Theorem A]. In particular, the representation $\rho^{\prime}$ is Aut-intrinsically Hodge-Tate but not crystalline. Thus, there is no inclusion relation in either direction between the classes of Aut-intrinsically Hodge-Tate representations and crystalline representations.
@@
2. On the non-normality of the field-theoretic subgroups of the outer automorphism groups of the absolute Galois groups of absolutely Galois MLFs
In the present §2, let $k$ be an MLF, $\overline{k}$ an algebraic closure of $k$, and $G$ a group of MLF-type. In the present §2, we discuss the outer automorphism groups of the absolute Galois groups of certain MLFs. In particular, we show that if $p_{k}$ is odd, $d_{k}$ is greater than one, and $k$ is an absolutely Galois MLF, i.e., the extension $k/k^{(d=1)}$ is a Galois extension, then the field-theoretic subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is not normal. This result generalizes certain theorems in [10] and [7].
@@
We shall refer to the image of the natural injective group homomorphism $\operatorname{Aut}(k)\hookrightarrow\operatorname{Out}(G_{k})$ [cf., e.g., [8], Proposition 2.1] as the field-theoretic subgroup of $\operatorname{Out}(G_{k})$.
@@
It is well-known that the isomorphism class of the field $k$ itself cannot be reconstructed group-theoretically from $G_{k}$ in general. For example, let $p$ be an odd prime number, $k_{0}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{Q}_{p}(\zeta_{p},\sqrt[p]{p})$ and $k_{1}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{Q}_{p}(\zeta_{p},\sqrt[p]{p+1})$, where $\zeta_{p}$ is a primitive $p$-th root of unity. Fix an algebraic closure $\overline{k_{0}}$ of $k_{0}$ and an algebraic closure $\overline{k_{1}}$ of $k_{1}$. Then it follows from the main theorem of [11] that the group $G_{k_{0}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Gal}(\overline{k_{0}}/k_{0})$ is isomorphic to the group $G_{k_{1}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Gal}(\overline{k_{1}}/k_{1})$, while the field $k_{0}$ is not isomorphic to the field $k_{1}$.
Furthermore, the following observation illustrates, in a stronger sense, that a group isomorphism $G_{k_{0}}\cong G_{k_{1}}$ is not compatible with their ring structures. Let $\varphi\colon G_{k_{1}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}G_{k_{0}}$ be an arbitrary group isomorphism, and let $\Phi\colon\operatorname{Out}(G_{k_{1}})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\operatorname{Out}(G_{k_{0}})$ denote the induced group isomorphism. Then it follows from [7], Corollary 8.6, that $\Phi(\operatorname{Gal}(k_{1}/\mathbb{Q}_{p}(\zeta_{p})))\not\subset N_{\operatorname{Out}(G_{k_{0}})}(\operatorname{Aut}(k_{0}))$. In particular, the subgroup $\operatorname{Aut}(k_{0})\subset\operatorname{Out}(G_{k_{0}})$ is not a normal subgroup of $\operatorname{Out}(G_{k_{0}})$.
@@
A group of MLF-type is an example of a slim profinite group [cf., e.g., [8], Proposition 1.8].
@@
Assertion (1) follows directly from Lemma 2.4, (1), and [8], Lemma 1.8. Assertion (2) follows directly from [8], Lemma 1.8. Assertion (3) follows directly from Lemma 2.4, (3), [8], Lemma 1.8, and the commutativity of the following diagram:
@@
Let $k(G_{k})$ be the topological field defined in [7], Definition 5.2 [Note that if $d_{k}=1$, then $d(G_{k})=1$ — cf. [8], Proposition 3.6]. By definition, the group $k(G_{k})^{\times}$ is the multiplicative group of the field $k(G_{k})$. Moreover, it follows immediately from [7], Theorem 5.4, that $k(G_{k})$ is canonically isomorphic to $k$ as a topological field. It follows from the functoriality of the mono-anabelian reconstruction algorithm that the following diagram commutes:
@@
This assertion follows immediately from Proposition 2.5, (3), Lemma 2.7, the commutative diagram appearing in the statement of [8], Lemma 1.7, $\rm(i\hskip-0.80002pti\hskip-0.80002pti)$, and the functoriality of the transfer map. ∎
The second main theorem of the present paper — namely, the non-normality of the field-theoretic subgroup of $\operatorname{Out}(G_{k})$ — follows from some results in §1 and the techniques developed in [10] under the assumptions that $d_{k}$ is even, $p_{k}$ is odd, and $k$ is absolutely Galois.
@@
Theorem 2.9 is a generalization of the main theorem of [10] and one of the main results of [7], which appeared in Remark 2.2. The method of the proof of Theorem 2.9 is essentially the same as the method of the proof that was applied in [10] and is substantially different from the approach taken in [7].
@@
Suppose that the subgroup $G_{k}\subset G_{k^{(d=1)}}$ is a characteristic open subgroup. For example, if $k/k^{(d=1)}$ is an abelian extension, then the subgroup $G_{k}\subset G_{k^{(d=1)}}$ is characteristic and open [cf. [7], Theorem 6.3, $\rm(\hskip 1.79993pti\hskip 1.79993pt)$]. Then, by Proposition 2.5, (2), (3), we obtain the following exact sequence of groups:
@@
From the existence of this exact sequence and the fact that $\operatorname{Out}(G_{k^{(d=1)}})\neq 1$ for odd residue characteristic $p_{k}$ [cf., e.g., the discussion given at the final portion of [15], Chapter VII, §5], we deduce that $\operatorname{Aut}(k)\subsetneq N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$. Put another way, there exists an element $\alpha\in N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$ such that $\alpha\notin\operatorname{Aut}(k)$. As discussed in the discussion following Remark 2.2, in general, the subgroup $(k^{(d=1)})^{\times}\subset k^{\times}$ cannot be reconstructed from the group $G_{k}$ in a purely group-theoretic and functorial manner. Put another way, the subgroup $(k^{(d=1)})^{\times}\subset k^{\times}$ depends on the ring structure of $k$ in an essential way. On the other hand, by Theorem 2.8, if $k$ is an absolutely Galois MLF and $\alpha\in N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$, then the induced isomorphism $\alpha^{\times}$ preserves the subgroup $(k^{(d=1)})^{\times}\subset k^{\times}$ and $\alpha^{\times}|_{(k^{(d=1)})^{\times}}=\operatorname{id}_{(k^{(d=1)})^{\times}}$. It follows from this discussion that every element $\alpha\in N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$, which is not necessarily contained in $\operatorname{Aut}(k)$, preserves certain structures of ring-theoretic nature.
@@
Let $\alpha$ be an element of $N_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k))$. Suppose that $k$ is an absolutely Galois MLF. As is shown in [10], Lemma 3.3, it holds that $\alpha_{+}|_{k^{(d=1)}_{+}}=\operatorname{id}_{k^{(d=1)}_{+}}$.
In what follows, we provide an alternative proof of this fact, which is stated in [10], by means of Theorem 2.8. What we need to show is that, for $\alpha\in\operatorname{Out}(G_{k})$, if $\alpha$ is $\mathbb{Q}_{p_{k}}^{\times}$-preserving, then it is also $(\mathbb{Q}_{p_{k}})_{+}$-preserving. Since $\alpha$ is $\mathbb{Q}_{p_{k}}^{\times}$-preserving, we have $\alpha^{\times}|_{(k^{(d=1)})^{\times}}=\operatorname{id}_{(k^{(d=1)})^{\times}}$. In particular, we have $\alpha^{\times}|_{\mathcal{O}^{\times}_{k^{(d=1)}}}=\operatorname{id}_{\mathcal{O}^{\times}_{k^{(d=1)}}}$. Therefore, on $2p_{k}\mathcal{I}_{k}$, we have $\alpha_{+}=\operatorname{id}_{2p_{k}\mathcal{I}_{k}}$. Since, for an arbitrary $x\in k_{+}$, there exists an integer $n$ such that $p_{k}^{n}x\in 2p_{k}\mathcal{I}_{k}$ [cf. [8], Lemma 1.2, $\rm(\hskip-0.59998ptv\hskip-0.80002pti)$], we conclude that $\alpha_{+}=\operatorname{id}_{k_{+}}$.
@@
Since $G$ is a topologically finitely generated profinite group [cf. Theorem 1.1], for any integer $n>0$, there exist only finitely many subgroups of $G$ of index $n$. In particular, the set of all characteristic subgroups of $G$ of finite index forms a fundamental system of open neighborhoods of $1\in G$. Hence, it follows from [18], Proposition 4.4.3, that $\operatorname{Aut}(G)$ and $\operatorname{Out}(G)$ have natural profinite group structures. In the remainder of the present paper, we endow $\operatorname{Aut}(G)$ and $\operatorname{Out}(G)$ with these profinite group structures.
Let us consider the group homomorphism $\operatorname{Out}(G)\to\operatorname{Aut}_{\mathbb{Q}_{p(G)}}(k_{+}(G))$ defined via the mono-anabelian reconstruction algorithm [cf. [8], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$, [10], Lemma 1.2]. We equip $\operatorname{Aut}_{\mathbb{Q}_{p(G)}}(k_{+}(G))$ with the topology induced by the $p(G)$-adic topology of $k_{+}(G)$, i.e., the relative topology with respect to
@@
By construction, the map $\operatorname{Aut}(G)\to\operatorname{Aut}(\mathcal{O}^{\times}(G))$ factors through the composite $\operatorname{Aut}(G)\to\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})\to\operatorname{Aut}(\mathcal{O}^{\times}(G))$, where we write $\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})$ for the subgroup of $\operatorname{Aut}(G^{\mathrm{ab}})$ consisting of elements that preserve the subgroup $\mathcal{O}^{\times}(G)\subset G^{\mathrm{ab}}$. We equip $\operatorname{Aut}(G^{\mathrm{ab}})$ with the profinite topology as in the case of $\operatorname{Aut}(G)$. The subgroup $\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})$ is endowed with the subspace topology induced from $\operatorname{Aut}(G^{\mathrm{ab}})$. It is clear that the map $\operatorname{Aut}(G)\to\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})$ is continuous. Moreover, since the profinite topology and the compact-open topology on $\mathrm{Aut}(G^{\mathrm{ab}})$ coincide [cf., e.g., [18], Theorem 4.4.2], the map $\operatorname{Aut}_{\mathcal{O}^{\times}(G)}(G^{\mathrm{ab}})\to\operatorname{Aut}(\mathcal{O}^{\times}(G))$, which is defined by $\varphi\mapsto\varphi|_{\mathcal{O}^{\times}(G)}$, is continuous. Thus, the action $\operatorname{Aut}(G)\curvearrowright\mathcal{O}^{\times}(G)$ is continuous. This completes the proof of Lemma 2.13. ∎
When $d_{k}>1$, let us consider the basis $y_{1},\ldots,y_{d_{k}}$ of $k_{+}$ over $\mathbb{Q}_{p_{k}}$ corresponding to the basis of $k_{+}(G_{k})$ that appeared in the discussion preceding Lemma 1.2 via the isomorphism of topological groups $k_{+}(G_{k})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}k_{+}$ of [8], Proposition 3.11, $\rm(i\hskip-0.80002ptv\hskip-0.59998pt)$. In the remainder of the present paper, we always equip $k_{+}$ with this basis, which allows us to identify $\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(k_{+})$ with $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$. We equip $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ with its natural $p_{k}$-adic Lie group structure.